Three boxes, each of mass 16 kg are on a frictionless table, connected by massless strings. A force of tension T1 pulls on the right most box (A) such that the three boxes accelerate in the positive horizontal direction at a rate of a 1.1 m/ 1) What is the magnitude of T1? 2) What is the net horizontal force on box A? N Submit 3) What is the net force that box B exerts on box A? N Submit 4) What is the net horizontal force on box B? N Submit
5) What is the net force box C exerts on box B? N Submit
Answer #1

a)

T1 = ( m1 + m2 + m3) a = 3 * 16 * 1.1 = 52.8 N

=======

b)

Considering mass B and C

T3 = m3 * a = 16 * 1.1 =17.6 N

Considering FBD of Mb

T2 = m2 a + T3

T2 = 16 * 1.1 + 17.6 = 35.2 N

net force on A

Fa = T1 - T2 = 17.6 N

======

c)

T2 = 35.2 N

======

d) Fb = 35.2 - 17.6 = 17.6 N

=====

e) T3 = 17.6 N

======

Comment in case any doubt, will reply for sure.. Goodluck